In a certain store, blue shirts are one-fifth the price of red shirts. If b blue shirts cost d dollars, what is the cost of r red shirts?
a) 5dr
b) 5brd
c) (rd)/(5b)
d) (5rd)/b
e) (5b/d)(r)

Found 2 solutions by math_tutor2020, Alan3354:
Answer by math_tutor2020(3817)


b = number of blue shirts
r = number of red shirts
d = cost of buying b blue shirts

b blue shirts cost d dollars
1 blue shirt costs d/b dollars. I divided both parts by b.

Example:
b = 4, d = 8
4 blue shirts cost 8 dollars
1 blue shirt costs 8/4 = 2 dollars


1 blue shirt is 1/5 the cost of 1 red shirt
Phrased in reverse, 1 red shirt is 5 times the cost of 1 blue shirt.

So,
cost of 1 red shirt = 5*(cost of 1 blue shirt)
cost of 1 red shirt = 5*(d/b)
cost of 1 red shirt = (5d)/b

The last step is to multiply that expression by r to find the cost of buying r number of red shirts.
We go from (5d)/b to (5rd)/b

Answer: (5rd)/b which is choice (d)


Answer by Alan3354(69443)